Net Present Value Method-Annuity Briggs Excavation Company is planning an investment of $109,400 for a bulldozer. The bulldozBriggs Excavation Equal Annual Net Cash Flow Cash inflows: Cash outflows:b. Determine the net present value of the investment, assuming that the desired rate of return is 15%. Use the present value
Answer #1

Solution:

Briggs Excavation Company

  1. Determination of equal annual net cash flows from operating the bulldozer:

Briggs Excavation

Equal annual net cash flow:

Cash Inflows:

Hours of operation

2,000

Revenue per hour

$115

Revenue per year

$230,000

Cash outflows:

Hours of operation

2,000

Fuel cost per hour

$50

labor cost per hour

$38

Total fuel and labor costs per year

$88 per hour

($176,000)

Annual maintenance cost

($20,000)

Annual net cash flow

$34,000

  1. Determination of the net present value of investment assuming desired rate of return as 15%:

Present value of annual net cash flows –

= annual net cash inflows x present value of annuity of $1 at 15% for 5 years

= $34,000 x (P/A, 15%, 5)

= 34,000 x 3.352 = $113,968

Less: initial investment $104,900

Net present value = 113,968 – 104,900 = $9,068

  1. Yes, Briggs Excavation should opt to invest in purchasing the bulldozer, as the investment earns a positive net present value at the company’s minimum desired rate of return of 15%.
  2. Determination of the operating hours so that the present value of cash outflows equal the amount to be investe

Amount to be invested = (P/A, 15%, 5) x [(number of hours x 115) – (number of hours x 88) – 20,000] = $109,400

$109,400 = 3.352 x [(number of hours x 115) – (number of hours x $88) – 20,000

109,400 = number of hours x $385.50 – number of hours x $295 – 67,040

$176,440 = number of hours x $90.50

Number of hours = 176,440/90.50 = 1,950           

Hence, the bulldozer’s operating hours must exceed 1,950 each year so as to justify the investment of $109,400.